An airplane starts from rest on the runway. The engines exert a constant force of 78.0 kN on the body of the plane (mass 9.20 × 104 kg) during takeoff. How far down the runway does the plane reach its takeoff speed of 74.7 m/s?
Answer #1

acceleration acts on plane

a = F/m = 78000 / ( 9.2* 10^4)

a = 0.8478 m/s^2

using 3rd equation of motion

v^2 = 2 a d

74.7^2 = 2* 0.8478* d

d = 74.7^2 / ( 2* 0.8478)

d = 5580.09 / 1.6956

d = 3290.82 m

=======

Comment before rate in case any doubt, will reply for sure.. goodluck